Can you help me find the factored form of +f+?
If +y+= f(x) is a cubic polynomial with the leading cofficent of +a+=+2+ and +f+(-2) = +f+(1) = +f+(2) = 0.

Answer by ikleyn(52779)

.
Can you help me find the factored form of +f+?
If +y+= f(x) is a cubic polynomial with the leading coefficent of +a+=+2+ and +f+(-2) = +f+(1) = +f+(2) = 0.
~~~~~~~~~~~~~~~~~~~~`

f(x) = 2*(x-(-2))*(x-1)*(x-2) = 2*(x+2)*(x-1)*(x-2).